# Tangling-Untangling Cycle for Efficient Learning

### v1

@@ IV Hierarchical Extension and Connection with Fractal Geometry
Figure 3: Hierarchical extension of TUC: a) spatial extension by Cartesian product of TUC, which forms the fundamental group of the circles; b) temporal extension by alternating between untangling and tangling in time domain.
Proposition 3. (Cartesian Product of TUC forms the Fundamental Group of the Circle).
@@
By recursively applying Cartesian products of TUC, we can construct a topological structure with increasing dimensionalities. Fig. 3a illustrates a toy example for such an extension of increasing dimensionalities (from a pair of circles to a torus). Analogous to the extension of circles to the fundamental group of circles, we can construct the fundamental group of TUC as $\textbf{S}^{1}\times...\times\textbf{S}^{1}$.
@@
An alternative hierarchical extension of TUC is based on the temporal depth [35] where the process of untangling and tangling occurs in an alternating manner. Temporal depth or hierarchy is formed based on allostasis [35], as shown in Fig 3b. This construction will be useful for characterizing the dynamics of TUC, such as error feedback and temporal stability, especially in the presence of heterogeneous interactions (e.g., sensory interactions with motor). The alternating between tangling and untangling operators is also analogous to the alternating between sleep and wake modes in sleep-wake cycles, as we will elaborate in the next section.
Proposition 4. (Allostatic concatenation of TUCs forms temporal hierarchy.
@@
Temporally hierarchical extension of TUCs can support continual learning by alternating between the processes of untangling and tangling in the time domain.
@@
A more subtle construction of hierarchical TUC is based on fractal geometry [28]. We must extend the strategy context dependency across different scales to support fractal TUC. This can be implemented by the positioning trick [26], which was inspired by Hippocampal indexing theory [46]. The indexing variables $I$ induce inter-scale context dependency - e.g., in structure-mapping theory for analogy [16], higher-order relations are modeled based on the relation between objects instead of attributes of objects (so-called systematicity principle). The implementation of such principle requires the context to be the object X instead of class label C, which requires a fractal design such as the six-layered organization of cortical column. Such hierarchical organization of structures into TUCs might explain the magical number of six (e.g., six degrees of separation) or seven [30].
Figure 4: Illustration of sleep-wake cycle (SWC): untangling occurs during the wake time (perturbation state) and tangling occurs during the sleep time (default mode).
V Biological Connections with PNG and SWC
@@
The biological implementation of the untangling operator is based on the connection of the polychronization neural groups (PNG) theory in [24] with the locally subspace untangling hypothesis in [9]. A direct consequence of direct-fit-to-nature is over-parameterization [20] to increase the predictive power without explicitly modeling the underlying generative structure of the world (i.e., intelligence without representation [3]). PNG offers a plausible computational model for supporting such mindless yet powerful generalizations. The number of PNG [24] can exponentially increase as more neurons are involved. Intuitively, one can argue that a high-level PNG with more neurons will have more representation power than a low-level PNG, which makes it a promising candidate for implementing the direct-fit-to-nature model [20].
@@
The combination of PNG with SWC [13] allows us to study cognitive functions from a novel TUC perspective. Using TUC as the vocabulary, we attempt to construct a computational model for SWC as the first step toward a more comprehensive neural syntax [4]. As application examples, we consider three tasks here: spatial navigation, object recognition, and motor control.
Untangling during the wake mode (perturbation state) is demonstrated by place cells with unidirectional place fields. The 1D travel sequence dictates the spatio-temporal context, which helps resolve the ambiguity in space [4]. Tangling during the sleep mode (default state) is demonstrated by omnidirectional place cells [34] that do not require a temporal context. Memory consolidation [44, 49] implements the aliasing or tangling by binding all PNG which shared the same context C together. The collapse of cloned variables occurs at the hippocampus where the encoding of spatio-temporal context (index to the local map or position on the global map) is handled by CA3-CA1 collateral system (pp. 326 [4]). During the consolidation of semantic memory (collection of CIR) from episodic memory (collection of CDR), the PNG-based time-locking mechanism eliminates the spatio-temporal context to support context-independent inference for generalization.
[…]
VI Computational Modeling of Sensorimotor and Social Interactions
TUC-based local computation for direct-fit inspired us to propose a new conceptual framework called “localized embodied cognition” with the locality principle in mind from problem formulation to solution algorithms. We still use the world “as its own model” [3], but the way local embodied cognition never attempts to solve a global problem locally. An organism needs to couple sensory with motor systems to achieve various goals but never stores knowledge about the global environment. Instead, the bodily interactions with the environment (situatedness) are persistently maintained by the neocortex (a single book)’s interaction with the hippocampus (a large library). The cognitive map in the hippocampus offers a situational context to interpret the sensorimotor interaction occurring in the neocortex. In this section, we study sensorimotor interaction and hippocampal-neocortical interaction using the newly-constructed neural syntax.
@@ VI Computational Modeling of Sensorimotor and Social Interactions
One of the biggest mysteries in studying cortical columns is that they are structures without a function [23]. Instead, several canonical microcircuits such as Douglas-Martin model [10] and Haeusler-Maass model [18] have been studied for predictive coding in the literature. These models have revealed a remarkable correspondence between the microcircuitry of the cortical column and the connectivity required by predictive coding. Cortical columns do not have a function because they work like swiss-army knivies. By connecting with different sensory and motor areas, they offer a universal processing algorithm [32]. TUC-based modeling allows us to gain a deeper understanding of sensorimotor interaction at the system level without the distraction from lower-level implementations.
@@
SMC-based Social Interaction
@@
. Carrying the “more is different” principle one step further, we can envision a bigger cycle connecting a pair of SMCs. When one person becomes another’s environment, the communication between two SMCs forms another bigger cycle known as non-verbal communication. The communication bandwidth of such non-verbal communication is limited by the locality of our sensory organs. Even though within each neocortex, the hierarchical organization maximizes the information flow inward, social communication (flowing outward) has to wait for evolution to improve its efficiency.
Figure 7: The cycle connecting speech production by motor cortex and language comprehension by auditory cortex has a dual structure to SMC because social communication between two people is connected by CIR (tangled states).
The fundamental role played by natural language is to facilitate communication among people - from non-verbal to verbal. Fig. 7 shows a modification of SMC to abstract the verbal communication between two people A and B, where motor and auditory cortex are responsible for speech generation and language comprehension, respectively. Note that unlike the original SMC in Fig. 6, the interface between two TUCs is in the tangled state (CIR). This difference has a fundamental impact on our internal (self-conscious) and external (social) cognitive function. For internal communication between sensory and motor cortex, the interface (inverse and forward in Fig. 6) connects two untangled states and there is no bottleneck in terms of information flow. For external (social) communication between two different person, the interface (inverse and forward in Fig. 7) connects two tangled states and faces an information flow bottleneck.
According to the simulation theory of mind-reading [15], mirror neurons can help us understand and empathize with the emotions and intentions behind others’ non-verbal cues. Modeling the new cycle in Fig. 7for understanding the social nature of language might offer new insight to the neural exploitation hypothesis [14]. According to this hypothesis, social cognition and language are both connected with action through the support of mirror neurons. However, the difference between non-verbal and verbal communication seems to be determined by their organizational principles. Non-verbal communication such as sign language does not admit hierarchical organization; while verbal communication such as Chomsky’s hierarchy [5] in natural language offers an evolutionary solution to maximize the information flow within the new cycle.
VII Conclusions and Future Works
In this paper, we have constructed a new TUC-based theory to solve the long-standing learning problem. Exploiting the blessing of dimensionality by inducing context dependency allows us to untangle any data manifold so linear classifiers will be sufficient. Then the tangling operator based on integral transform collapses the induced context variables for generalization. Such mapping between tangled CIR and untangled CDR can be extended into hierarchical representations by Cartesian product and fractal geometry. The computational efficiency of TUC-based learning algorithms is supported by the biological implementation by SWC. Sleep and wake modes in SWC elegantly correspond to tangling and untangling operators due to the absence of sensory inputs during the sleep mode. SWC allows us to construct a neural syntax for an improved understanding of cognitive functions associated with sensorimotor interaction and social interaction.

### v2

@@ 4 Hierarchical Extension: from Cycles to Torus and Chain
Figure 4: Hierarchical extension of SWC: a) spatial extension by Cartesian product of SWC, which forms the fundamental group of the circles; b) temporal extension by alternating between untangling and tangling in time domain.
Proposition 3. (Cartesian Product of DEC forms the Fundamental Group of the Circle).
@@
By recursively applying Cartesian products of DEC, we can construct a topological structure with increasing dimensionalities. Fig. 4a illustrates a toy example for such an extension of increasing dimensionalities (from a pair of circles to a torus). Analogous to the extension of circles to the fundamental group of circles, we can construct the fundamental group of DEC as $\textbf{S}^{1}\times...\times\textbf{S}^{1}$.
@@
An alternative hierarchical extension of DEC is based on the temporal depth [56] where the process of untangling and tangling occurs in an alternating manner. Temporal depth or hierarchy is formed based on allostasis [56], as shown in Fig. 4b. This construction will be useful for characterizing the dynamics of DEC, such as error feedback and temporal stability, especially in the presence of heterogeneous interactions (e.g., sensory interactions with motor). The alternating between tangling and untangling operators is also analogous to the alternating between sleep and wake modes in sleep-wake cycles, as we will elaborate in the next section.
Proposition 4. (Allostatic concatenation of DECs forms temporal hierarchy.
@@
Temporally hierarchical extension of DECs can support continual learning by alternating between the processes of untangling and tangling in the time domain.
@@
A more subtle construction of hierarchical DEC is based on fractal geometry [44]. We must extend the strategy context dependency across different scales to support fractal DEC. The positioning trick can implement this [43], which was inspired by Hippocampal indexing theory [70]. The indexing variables $I$ induce inter-scale context dependency - e.g., in structure-mapping theory for analogy [26], higher-order relations are modeled based on the relation between objects instead of attributes of objects (so-called systematicity principle). The implementation of such principle requires the context to be the object X instead of class label C, which requires a fractal design such as the six-layered organization of cortical column. Such hierarchical organization of structures into DECs might explain the magical number of six (e.g., six degrees of separation) or seven [47].
5 Biological Implementations and Applications
@@
The biological implementation of both entanglment and disentanglement operators is based on the polychronization neural groups (PNG) theory [36]. A direct consequence of direct-fit-to-nature is over-parameterization [30] to increase the predictive power without explicitly modeling the underlying generative structure of the world. PNG offers a plausible computational model for supporting such mindless yet powerful generalizations. The number of PNG [36] can exponentially increase as more neurons are involved. Intuitively, one can argue that a high-level PNG with more neurons will have more representation power than a low-level PNG, which makes it a promising candidate for implementing the over-parameterized model required by direc-fit [30].
@@
The connection of PNG with SWC [21] allows us to study cognitive functions from a novel DEC perspective. Using DEC as the vocabulary, we attempt to construct a computational model for SWC as the first step toward a more comprehensive neural syntax [9]. As application examples, we consider three tasks here: spatial navigation, object recognition, and motor control.
Untangling during the wake mode (perturbation state) is demonstrated by place cells with unidirectional place fields. The 1D travel sequence dictates the spatio-temporal context C, which helps resolve the ambiguity in space [9]. Tangling during the sleep mode (default state) is demonstrated by omnidirectional place cells X [55] that do not require a temporal context C. Memory consolidation [69, 74] implements the aliasing or tangling by binding all PNG which share the same context C together. The collapse of cloned variables occurs at the hippocampus where the encoding of spatio-temporal context (index to the local map or position on the global map) is handled by CA3-CA1 collateral system (pp. 326 [9]). During the consolidation of semantic memory (collection of CIR) from episodic memory (collection of CDR), the PNG-based time-locking mechanism eliminates the spatio-temporal context to support context-independent inference for generalization.
@@
6 Computational Modeling of Sensorimotor and Social Interactions
DEC-based local computation for direct-fit inspired us to propose a new conceptual framework called “localized embodied cognition” with the locality principle in mind from problem formulation to solution algorithms. We still use the world “as its own model” [5], but the way local embodied cognition never attempts to solve a global problem locally. An organism needs to couple sensory with motor systems to achieve various goals but never stores knowledge about the global environment. Instead, the bodily interactions with the environment (situatedness) are persistently maintained by the neocortex (a single book)’s interaction with the hippocampus (a large library). The cognitive map in the hippocampus offers a situational context to interpret the sensorimotor interaction occurring in the neocortex. In this section, we study sensorimotor interaction and hippocampal-neocortical interaction using the newly-constructed neural syntax.
@@ 6 Computational Modeling of Sensorimotor and Social Interactions
One of the biggest mysteries in studying cortical columns is that they are structures without a function [35]. Instead, several canonical microcircuits such as Douglas-Martin model [18] and Haeusler-Maass model [28] have been studied for predictive coding in the literature. These models have revealed a remarkable correspondence between the microcircuitry of the cortical column and the connectivity required by predictive coding. Cortical columns do not have a function because they work like swiss-army knivies. By connecting with different sensory and motor areas, they offer a universal processing algorithm [53]. DEC-based modeling allows us to gain a deeper understanding of sensorimotor interaction at the system level without the distraction from lower-level implementations.
@@
SMC-based Social Interaction.
@@
Carrying the “more is different” principle one step further, we can envision a bigger cycle connecting a pair of SMCs. When one person becomes another’s environment, the communication between two SMCs forms another bigger cycle known as non-verbal communication. The communication bandwidth of such non-verbal communication is limited by the locality of our sensory organs. Even though within each neocortex, the hierarchical organization maximizes the information flow inward, social communication (flowing outward) has to wait for evolution to improve its efficiency.
The fundamental role played by natural language is to facilitate communication among people - from non-verbal to verbal. If one concatenates the SMCs of two people (for communication purposes), we observe a larger cycle $SMC_{A}\leftrightarrows SMC_{B}$. According to the simulation theory of mind-reading [24], mirror neurons can help us understand and empathize with the emotions and intentions behind others’ nonverbal cues. Modeling the bigger cycle (refer to Appendix A for a more detailed discussion about the ecological origin of languages) for understanding the social nature of language might offer new insight to the neural exploitation hypothesis [23]. Since nature is good at recycling its tricks, the originally evolved sensorimotor integration as shown in Fig. 6 might be adapted for the new function of language as a more efficient alternative to nonverbal communication.
7 Conclusions, Limitations, and Broader Impact
Conclusions: This paper studies the computational modeling of sleep-wake cycle (SWC) from a direct-fit perspective. The wake mode entangles the sensory observations by binding them into gnostic cells. The sleep mode may consolidate memories by replaying the waking experience. Memory consolidation is abstracted by an integral transform in this paper, which disentangles the manifold for generalization. The pair of entanglement and disentanglement operators implement the mapping between CIR and CDR, demonstrating the importance of context dependency to learning and memory. The hierarchical extension of disentanglement-entanglement cycles into high-order topological structures (e.g., torus and chain) is an attempt to enhance the computational modeling capability of SWC for more sophisticated cognitive tasks.
Limitations: Our paper is inspired by many cycles discovered by Buzsaki in the rhythms of the brain [9]. Its biggest limitation is the lack of experimental verification. As pointed out in [9], studying the brain’s default mode is even more challenging than the perturbation mode. Experiments related to sleep-dependent memory consolidation [69] only indirectly provide evidence to support the role of SWC in semantic memory. Many cycles discovered in this paper are limited to theoretical appeal, whose practical implications and connections remain to be established. Due to space limitation, we have skipped the connection of DEC with limit-cycle behavior of Wilson-Cowan model [76] and its related hysteresis cycle (the physical origin of memory).
Broader Impacts: Understanding the sleep-wake cycle paves the path to unlocking the secret of learning and memory. At the device level of individual neurons, limit-cycle behavior has been well studied by the Wilson-Cowan model [76]. At the circuit level of neural groups, canonical circuits [39] have formed the cycle for predictive processing. This paper marks the first step toward discovering cycles at the system level of neural areas. Since SWC is fundamental to almost all cognitive functions, its computational modeling might facilitate the study of both normal behaviors and neurological disorders such as autism and anxiety. Studying the default state of the brain (sleep mode) is deemed to complement the study of the perturbation state (wake mode).
Appendix A Ecological View of Sleep/Wake Modes and Sleep-Wake Cycle
The Brain’s Default State: Sleep Mode.
According to Gyorgy Buzsaki [9], rest and sleep are characterized by self-organized oscillations in the absence of environmental inputs. Without an outside influence, the evolving state in the sleep mode is deterministic (i.e., completely predictable). The role of sleep in memory consolidation has been observed in the experiments with sleep spindles [64]. It was hypothesized that during spindles, specific neuronal populations become more strongly modulated through virtual replay [37]. Experiments with the visual cortex and hippocampus have suggested that they support simultaneous reactivation of coherent memory traces during slow-wave sleep (SWS) and rapid-eye-movement (REM) sleep [37]. All these experimental findings are consistent with Proposition 2 in our SWC theory. In summary, Memory consolidation is a process by which the hippocampus guides the organization of the information stored in the neocortex [68]. The consequence of self-organization is that the consolidated memory becomes independent of the hippocampus - e.g., from episodic memory (context-dependent) to semantic memory (context-independent).
The Brain’s Perturbation State: Wake Mode.
According to Karl Friston, “brain dynamics constantly shift from the complex to the predictable” (pp. 111, [9]). The waking brain has been hypothesized to be optimized for encoding memories [60]. From sensory perception to motor control, goal-oriented behaviors offer a unifying perspective of understanding the brain’s state during the wake mode. According to Hebbian postulate [32], the synchronization of cell assemblies (a.k.a., polychronization of neural groups [36]) drives the brain dynamics “from stable incoherence through dynamic instability to complete entrainment” [20]. Unlike the deterministic default mode, the stochastic nature of environmental inputs dictates the brain to strike a good balance between specification and generalization. By encoding history into memory, complex systems manage to make reliable predictions. Proposition 1 in our SWC theory is consistent with this line of reasoning because CDR facilitates the prediction by inductive bias.
Sleep-Wake Cycle and Memory Consolidation.
Not only humans but also mammals go to sleep [66]. Across mammals, the amount and nature of sleep are often correlated with their body size, age, and ecological variables (e.g., terrestrial or aquatic). The role of the SWC in energy conservation and nervous system recuperation has been widely observed in mammals. The recognition of SWC as a biological rhythm controlled by brain oscillators (i.e., beyond the framework of homeostatic principles) has only been more widely accepted since the 1960s [42]. The basis of SWC has been related to the process of toxins’ accumulation and removal [42]. The neurobiology of SWC was reviewed in [21], highlighting the sleep architecture, circadian regulation, and regulatory feedback. The close relationship between memory encoding/consolidation and SWC was studied by Buzsaki in 1998 [7] (a more recent review can be found in [17]). The integral transform in Proposition 2 might be a mathematical abstraction of synaptic modifications in the hippocampus during sleep. The importance of sleep patterns to the preservation of experience-induced synaptic modifications has been documented by various studied on sleep deprivation [40].
Appendix B On Ecological Origin of Languages
The origin of language was once a taboo among linguists due to a lack of scientific methodology [34]. In 1950, Alan Turing invented an imitation game (later known as the Turing test) for artificial intelligence (AI) [72]. In the Turing test, a human evaluator would judge natural language conversations between a human and a machine designed to generate human-like responses. Recent breakthroughs in generative AI (e.g., ChatGPT developed by OpenAI [6]) have shown convincing evidence for machine’s nearly passing the Turing test (“Essentially, whether I pass the Turing Test might depend on the context and the criteria set by the human interacting with me.” - quoted from ChatGPT). However, passing a Turing test does not imply the possession of minds like humans do, as demonstrated by John Seale in Chinese room thought experiments in 1980. The fundamental reason behind the Chinese room argument is that the Turing test is a top-down approach to understanding intelligence [9].
An alternative approach is bottom-up (a.k.a. inside-out), which requires a deeper understanding of the mechanism underlying the structure and function of neural systems. As discovered by Vernon Moutcastle in 1957 [52], there is no visible difference among the six-layer organization of the neocortex no matter which area we look into. Based on this observation, he hypothesized that all cognitive functions, from spatial navigation and motor control to vision and language, result from a universal cortical processing algorithm. According to Theodosius Dobzhansky, “nothing in biology makes sense except in the light of evolution.” The evolution of mammalian brains offers fruitful insight into the functional organization of the neocortex. As Hans Moravec says in [49], abstract thought “is a new trick, perhaps less than 100 thousand years old….effective only because it is supported by this much older and much more powerful, though usually unconscious, sensorimotor knowledge.” To understand language, we must understand sensorimotor learning first.
Following the clue offered by the evolution of mammalian brains, one can understand the allocortex that is sandwiched between the old and new brain for spatial navigation [71]. The discovery of place cells in the hippocampus in 1978 [55] marked a milestone in our journey to a better understanding of how the brain works. This line of research has accumulated many more discoveries (e.g., grid cells [50, 51]) and computational models related to navigation, contributing to our knowledge about how to build cognitive maps [75]. Along this line of reasoning, the relationship between place cells and memory consolidation [69] is particularly fascinating because if Mountcastle were correct, the trick used by nature to create gnostic units such as place cells is likely to be recycled for other cognitive functions (e.g., concept cells in object recognition [59, 58]).
To solve the navigation problem, context (for specification) and aliasing (for generalization) are two clever tricks invented by nature. As shown in [9] (pp. 327), oscillatory timing of cell assemblies defines the spatiotemporal context for a currently coded item (e.g., the current position of the animal). During the wake mode (perturbation state), animals generate unidirectional place fields as context-dependent episodic memories about one-dimensional travel routes. During the sleep mode (default state), memory consolidation uses an aliasing trick to collapse unidirectional place cells with identical positions but different contexts into omnidirectional place cells. Unlike unidirectional place cells that are context-dependent, omnidirectional place cells are context-independent, implying better generalization. Such an aliasing trick becomes Maslow’s hammer [45] to tackle other cognitive tasks.
a)                                                         b)
Figure A1: Ecological origin of languages. a) language is needed to support the external sensorimotor cycle where the sensory cortex and motor cortex cannot directly communicate with each other but have to overcome the information bottleneck of the auditory cortex (language comprehension is the inverse pass transforming auditory-CIR into motor-CIR to facilitate motor control for speech generation); b) Chomsky hierarchy mirrors the nested organization of the physical world so the neocortex can recycle Maslow’s hammer in verbal communication.
Nature has cleverly used the aliasing trick, based on Maslow’s hammer, to solve the learning problem by memory. In simple terms, the spatio-temporal context in navigation becomes a response vector to a feature input in perception or action. The problem of sensorimotor learning is then transformed into a generalized navigation problem of memorizing a “travel sequence” in the latent space [67] that comprises the concatenation of feature input and response vectors. During the wake mode, the brain encodes many such travel sequences, which are known to suffer from overfitting. However, during sleep mode, memory consolidation collapses episodic memory into semantic memory. This is achieved by binding polychronization neural groups with the same feature input but different responses (contextual variables) together for generalization. During sleep-wake cycles [74, 69], overfitted context-dependent representations (CDR) are self-organized into context-independent representations (CIR) without the bias-variance barrier [25] or the curse of dimensionality [4] because the brain does not solve any optimization problem but count on memory (Maslow’s Hammer) for adaptive behaviors.
The evolution of sensorimotor learning greatly contributes to human flexible behaviors in terms of perception and action [67]. However, social communication between two people requires another clever recycling of Maslow’s Hammer. Unlike abstract thoughts that can be done internally in our mind, social communication requires the completion of an external cycle connecting the sensorimotor systems of two people. Such an external loop is at the mercy of the information bottleneck of our sensory organs because we cannot directly communicate with others like a three-body civilization (in SCIFI fiction “Three-Body”) that cannot lie. Instead, we have to count on external communication channels to exchange our thoughts. Visual communication such as sign language is limited by its capacity; natural language is more effective because the neocortex can recycle its sensorimotor skills learned from interacting with the world (corresponding to the internal cycle, as shown Fig. 6) to interacting with the peer (corresponding to the external cycle, as shown Fig. A1a). Such ecological explanation for the origin of natural languages is another evidence supporting Maslow’s Hammer used by evolution.
This line of reasoning explains the hierarchical organization discovered by Noam Chomsky in 1957 [10] and supports the Sapir–Whorf hypothesis [38] in linguistic relativity. The Sapir–Whorf hypothesis, also known as linguistic relativity, is a concept in linguistics that suggests the structure of a language affects its speakers’ worldview or cognition. In summary, Chomsky hierarchy (Fig. A1b) mirrors the nested organization of the physical world.